A wire has a mass per unit length of 10 g/m. What
Answer #1

Force in magnetic field is given by

F = i*B*L

Force due to gravity

Fg = m*g

mg = iBL

i - mg/BL

m/L = 10 gm/m = 0.01 kg/m

g = 9.8 m/sec^2

B = 0.1 T

So,

i = 0.01*9.8/0.1 = 0.98 Amp.

Correct option is 1